Peanuts worth $2.50 a pound were mixed with cashews worth $3.25 a pound. How many pounds of each were used if there were ten more pounds of peanuts that cashews and if the value of the mixture was $122.75?
Answer by checkley77(12844)

2.50(x+10)+3.25x=122.75
2.50x+25+3.25x=122.75
5.75x=122.75-25
5.75x=97.75
x=97.75/5.75
x=17 pounds of cashews were used.
17+10=27 ponds of peanuts were used.
Proof;
2.50*27+3.25*17=122.75
67.50+55.25=122.75
122.75=122.75